### v2

# Tony Lindeberg

Abstract
This paper presents results of combining (i) theoretical analysis regarding connections between the orientation selectivity and the elongation of receptive fields for the affine Gaussian derivative model with (ii) biological measurements of orientation selectivity in the primary visual cortex, to investigate if (iii) the receptive fields can be regarded as spanning a variability in the degree of elongation.
From an in-depth theoretical analysis of idealized models for the receptive fields of simple cells and complex cells in the primary visual cortex, we have established that the directional selectivity becomes more narrow with increasing elongation of the receptive fields. By comparison with previously established biological results, concerning broad vs. sharp orientation tuning of visual neurons in the primary visual cortex, we demonstrate that those underlying theoretical predictions, in combination with these biological results, are consistent with a previously formulated biological hypothesis, stating that the biological receptive field shapes should span the degrees of freedom in affine image transformations, to support affine covariance over the population of receptive fields in the primary visual cortex.
Based on this possible indirect support for the working hypothesis concerning affine covariance, we formulate a set of testable predictions that could be used to, with neurophysiological experiments, judge if the receptive fields in the primary visual cortex of higher mammals could be regarded as spanning a variability over the eccentricity or the elongation of the receptive fields, and, if so, then also characterize if such a variability would, in a structured way, be related to the pinwheel structure in the visual cortex.
Keywords:
Receptive field Elongation Affine covariance Orientation selectivity Gaussian derivative Quasi quadrature Simple cell Complex cell Pinwheel Vision Theoretical neuroscience
journal: arXiv preprint
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Variability in image structures generated by viewing the same surface pattern for different slant angles, that is different angles between the surface normal and the viewing direction. Observe how the resulting perspective effects lead to perspective foreshortening in the vertical direction. (When taking these images, the distance between the object and the observer as well as the amount of zoom for the zoom lens was, however, also adjusted to keep the vertical size of the surface pattern roughly the same in the image domain.)
$\kappa=1$ $\kappa=\sqrt{2}$ $\kappa=2$ $\kappa=2\sqrt{2}$ $\kappa=4$ $\kappa=4\sqrt{2}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Variability in the elongation of affine Gaussian derivative receptive fields (for image orientation $\varphi=\pi/2$), with the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ increasing from $1$ to $4\sqrt{2}$ according to a logarithmic distribution, from left to right, with the larger of these scale parameters $\sigma_{2}$ kept constant. (top row) First-order directional derivatives of affine Gaussian kernels according to (1) for $m=1$. (bottom row) Second-order directional derivatives of affine Gaussian kernels according to (1) for $m=2$. (Horizontal axes: image coordinate $x_{1}$. Vertical axes: image coordinate $x_{2}$.)
1 Introduction
When observing objects and events in our natural environment, the image structures in the visual stimuli will be subject to a substantial variability, caused by the natural image transformations. Specifically, if observing a smooth local surface patch from different viewing directions and viewing distances, this variability can to first order of approximation be approximated by local affine transformations (the derivative of the projective mappings between the different views). Within the 4-D variability of general centered 2-D affine transformations, a 1-D variability in the slant angle of the surface normal relative to the viewing direction does, in terms of covariance properties, correspond to a variability in the elongation of the receptive fields, if we would like the responses to be possible to perfectly match under such variabilities (see Figures 1 and 2 for illustrations).
In (Lindeberg 2021; 2023) we have outlined a framework for how covariance properties with respect to geometric image transformations may constitute a fundamental constraint for the receptive fields in the primary visual cortex of higher mammals, to enable the visual computations to be robust under the variabilities in the image structures generated by the natural image transformations. According to the presented theory, based on axiomatically determined receptive field shapes derived from symmetry properties that reflect structural properties of the environment, in combination with additional constraints to guarantee consistency between image representations over multiple spatial and temporal scales, the population of receptive fields in the primary visual cortex ought to, according to this theory, obey covariance properties with respect to spatial affine transformations and Galilean transformations.
While overall qualitative comparisons between predictions from this principled theory have been successfully made to neurophysiological recordings of receptive fields of simple cells by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008), the publicly available data regarding full receptive field recordings are quite limited, why further experimental evidence would be needed to firmly either reject or support the stated hypotheses about affine covariance and Galilean covariance.
In the lack of such neurophysiological data regarding full receptive field recordings, one could, however, aim to instead obtain indirect cues regarding a possible variability in the degree of elongation of the receptive fields, by making use of the recordings of the orientation selectivity of visual neurons by Nauhaus et al. (2008), which show a substantial variability regarding broad vs. sharp tuning of the receptive fields in the primary visual cortex.
In a companion paper (Lindeberg 2024), we have established a strong direct link between the orientation selectivity and the elongation of the receptive fields according to the idealized generalized Gaussian derivative model for visual receptive fields (as will be summarized in Section 2). If we would assume that that model would constitute a sufficiently valid model for the population of simple and complex cells in the primary visual cortex, then we could by logical inference infer possible indirect support for the working hypothesis, in that the variability in the orientation selectivity of the receptive fields would correspond to a variability in the degree of elongation of the receptive fields.
1.1 The hypothesis about affine covariant receptive fields
If we assume that the visual system should implement affine covariant receptive fields, to make it possible for the visual system to better estimate shape properties of surfaces in the world, such as local surface orientation, then the property of affine covariance would make it possible to compute better estimates of local surface orientation, compared to a visual system that does not implement affine covariance, or a sufficiently good approximation thereof.
A general motivation for the wider underlying working hypothesis about affine covariance is that, if the population of receptive fields would support affine covariance in the primary visual cortex, or sufficiently good approximations thereof, then such an ability would support the possibility of computing affine invariant image representations at higher levels in the visual hierarchy (Lindeberg 2013b), or more realistically sufficiently good approximations thereof, over restricted subspaces or subdomains of the most general forms of full variability under spatial affine transformations of the visual stimuli.
Fundamentally, we cannot expect that visual perception should be able to implement full affine invariance. For example, from the well-known experience, that it is much harder to read a text upside-down, it is clear that visual perception cannot be regarded as invariant to spatial rotations in the image domain. From the expansion of the orientations of visual receptive fields according to the pinwheel structure of higher mammals, we can, however, regard the population of receptive fields as supporting local rotational covariance. If we look at a slanted surface in the world, we can get a robust perception of its surface texture under substantial variations of the slant angle, thus showing a robustness of visual perception under stretchings of visual image patterns that correspond to non-uniform scaling transformations (the perspective effects on a slanted surface patch can to first-order of approximation be modelled as a stretching of the image pattern along the tilt direction in image space, complemented with a uniform scaling transformation). If the visual receptive fields would span a variability under such spatial stretching transformations, then such a variability would precisely correspond to a variability in the anisotropy, or the degree of elongation, of the receptive fields.
1.2 Variability over the elongation of receptive fields
The overall theme of this article is thus to, based on a theoretical analysis of a relationship between the orientation selectivity and the degree of anisotropy or degree of elongation of receptive fields in the primary visual cortex, in combination with existing neurophysiological results concerning variabilities in the orientation selectivity of the receptive fields in relation to the pinwheel structure of higher mammals, address the question of whether the receptive fields in the primary visual cortex could be regarding as spanning a variability over the elongation of the receptive fields, to, in turn, support covariant image measurements under such variabilities, or, at least, sufficiently good approximations thereof.
If we, in particular, could assume that the spatial components of the receptive fields could be well modelled by affine Gaussian derivatives, then the property of affine covariance implies that there should be visual receptive fields for different degrees of anisotropy present in the visual system. If we further combine the theoretical results used in this paper, which will state that the degree of orientation selectivity is strongly dependent on the anisotropy or the elongation of the receptive fields, with the biological results established by Nauhaus et al. (2008), which show that the degree of orientation selectivity for neurons in the primary visual cortex varies both strongly and in relationship with the position on the cortical surface in relation to the pinwheel structure, then these results together are fully consistent with the hypothesis that the visual receptive fields in the primary visual cortex should span a variability in their anisotropy, thus consistent with the hypothesis that the receptive fields in the pinwheel structure should span at least one more degree of freedom in the affine group, beyond mere rotations (as already established in previous neurophysiological measurements regarding the pinwheel structure of the oriented receptive fields in the primary visual cortex of higher mammals).
We will also, more generally, use predictions from the presented theoretical analysis to formulate a set of explicit testable hypotheses, that could be either verified or rejected in future neurophysiological experiments. Additionally, we will formulate a set of quantitive measurements to be made, to characterize a possible variability in the anisotropy or elongation of receptive fields in the primary visual cortex, with special emphasis on the relationships between a possibly predicted variability in receptive field elongation and the pinwheel structure in the primary visual cortex of higher mammals.
2 Connections between the orientation selectivity and the elongation of receptive fields for the generalized Gaussian derivative model for visual receptive fields
For modelling the receptive fields in the primary visual cortex, we will make use of the generalized Gaussian derivative model for receptive fields.
2.1 Idealized models for simple cells
The purely spatial component of the receptive fields for the simple cells modelled as (Lindeberg 2021 Equation (23); see Figure 7 in that reference for illustrations)
$T_{\scriptsize\mbox{simple}}(x_{1},x_{2};\;\sigma_{\varphi},\varphi,\Sigma_{\varphi},m)\\ =T_{\varphi^{m},\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\sigma_{\varphi}^{m}\,\partial_{\varphi}^{m}\left(g(x_{1},x_{2};\;\Sigma_{\varphi})\right),$ (1)
and with joint spatio-temporal receptive fields of the simple cells according to (Lindeberg 2021 Equation (25); see Figures 10-11 in that reference for illustrations)
$\begin{split}T_{\scriptsize\mbox{simple}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},\varphi,v,\Sigma_{\varphi},m,n)\end{split}$
$\begin{split}&=T_{{\varphi}^{m},{\bar{t}}^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})\end{split}$
$\begin{split}&=\sigma_{\varphi}^{m}\,\sigma_{t}^{n}\,\partial_{\varphi}^{m}\,\partial_{\bar{t}}^{n}\left(g(x_{1}-v_{1}t,x_{2}-v_{2}t;\;\Sigma_{\varphi})\,h(t;\;\sigma_{t})\right),\end{split}$ (2)
where
  • $\varphi$ is the preferred orientation of the receptive field,
  • $\sigma_{\varphi}$ is the amount of spatial smoothing,
  • $\partial_{\varphi}^{m}=(\cos\varphi\,\partial_{x_{1}}+\sin\varphi\,\partial_{x_{2}})^{m}$ is an $m$:th-order directional derivative operator, in the direction $\varphi$,
  • $\Sigma_{\varphi}$ is a symmetric positive definite covariance matrix, with one of its eigenvectors in the direction of $\varphi$,
  • $g(x;\;\Sigma_{\varphi})$ is a 2-D affine Gaussian kernel with its shape determined by the covariance matrix $\Sigma_{\varphi}$
    $g(x;\;\Sigma_{\varphi})=\frac{1}{2\pi\sqrt{\det\Sigma_{\varphi}}}e^{-x^{T}\Sigma_{\varphi}^{-1}x/2}$ (3)
    for $x=(x_{1},x_{2})^{T}$,
  • $\sigma_{t}$ is the amount of temporal smoothing,
  • $v=(v_{1},v_{2})^{T}$ is a local motion vector, in the direction $\varphi$ of the spatial orientation of the receptive field,
  • $\partial_{\bar{t}}^{n}=(\partial_{t}+v_{1}\,\partial_{x_{1}}+v_{2}\,\partial_{x_{2}})^{n}$ is an $n$:th-order velocity-adapted temporal derivative operator,
  • $h(t;\;\sigma_{t})$ is a temporal Gaussian kernel with standard deviation $\sigma_{t}$.
This model builds upon the regular Gaussian derivative model for purely spatial receptive fields proposed by Koenderink and van Doorn (1984; 1987; 1992) and previously used for modelling biological fields by Young and his co-workers 1987; 2001; 2001, however, here also generalized to affine covariance according to Lindeberg (2013a; 2021).
2.2 Idealized models for complex cells
To model complex cells with a purely spatial dependency, we will use a quasi quadrature measure of the form (Lindeberg 2020 Equation (39))
${\cal Q}_{\varphi,\scriptsize\mbox{spat},\scriptsize\mbox{norm}}L=\sqrt{L_{\varphi,\scriptsize\mbox{norm}}^{2}+C_{\varphi}\,L_{\varphi\varphi,\scriptsize\mbox{norm}}^{2}},$ (4)
where
  • $L_{\varphi,\scriptsize\mbox{norm}}$ and $L_{\varphi\varphi,\scriptsize\mbox{norm}}$ constitute the results of convolving the input image with scale-normalized directional affine Gaussian derivative operators of orders 1 and 2:
    $L_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\\ =T_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (5)
    $L_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\\ =T_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (6)
  • $C_{\varphi}>0$ is a weighting factor between first and second-order information.
This model constitutes an affine Gaussian derivative analogue of the energy model of complex cells developed by Adelson and Bergen (1985) and Heeger (1992), and is consistent with the observation that receptive fields analogous to first- vs. second-order derivatives occur in pairs in biological vision (De Valois et al. 2000), with close analogies to quadrature pairs, as defined in terms of a Hilbert transform (Bracewell 1999, pp. 267–272).
Complex cells with a joint spatio-temporal dependency will, in turn, be modelled as
$({\cal Q}_{\varphi,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)=\sqrt{L_{\varphi,\scriptsize\mbox{norm}}^{2}+\,C_{\varphi}\,L_{\varphi\varphi,\scriptsize\mbox{norm}}^{2}},$ (7)
where
$L_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})=\\ =T_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})*f(x_{1},x_{2},t),$ (8)
$L_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})=\\ =T_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})*f(x_{1},x_{2},t),$ (9)
with the underlying space-time separable spatio-temporal receptive fields $T_{\varphi^{m},t^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})$ according to (2.1) for $n=0$.
Figure 3: Schematic illustration of the sine wave probe used for defining the orientation selectivity curve, by using a receptive field model with the fixed preferred orientation $\varphi=0$, and then exposing the receptive field to sine waves for different inclination angles $\theta$. (Horizontal axis: spatial coordinate $x_{1}$. Vertical axis: spatial coordinate $x_{2}$.)
First-order simple cell Second-order simple cell Complex cell
$\kappa=1$ Refer to caption Refer to caption Refer to caption
$\kappa=2$ Refer to caption Refer to caption Refer to caption
$\kappa=4$ Refer to caption Refer to caption Refer to caption
$\kappa=8$ Refer to caption Refer to caption Refer to caption
Figure 4: Graphs of the orientation selectivity for the idealized models of (left column) simple cells in terms of first-order directional derivatives of affine Gaussian kernels, (middle column) simple cells in terms of second-order directional derivatives of affine Gaussian kernels and (right column) complex cells in terms of directional quasi quadrature measures that combine the first- and second-order simple cell responses in a Euclidean way for $C_{\varphi}=C_{t}=1/\sqrt{2}$, and shown for different values of the ratio $\kappa$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the degree of orientation selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.)
2.3 Orientation selectivity curves for the idealized receptive field models
In (Lindeberg 2024) the responses of the above purely spatial models of receptive fields are calculated with respect to a static sine wave of the form (see Figure 3)
$f(x_{1},x_{2})=\sin\left(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta\right).$ (10)
Additionally, the responses of the above joint spatio-temporal models of receptive fields are calculated with respect to a moving sine wave of the form
$f(x_{1},x_{2},t)=\\ =\sin\left(\omega\cos(\theta)\,(x_{1}-u_{1}t)+\omega\sin(\theta)\,(x_{2}-u_{2}t)+\beta\right),$ (11)
with the velocity vector $(u_{1},u_{2})^{T}$ parallel to the inclination angle $\theta$ of the grating such that $(u_{1},u_{2})^{T}=(u\cos\theta,u\sin\theta)^{T}$.
In summary, the theoretical analysis in (Lindeberg 2024) shows that the resulting orientation selectivity curves for the first-order simple cells, the second-order simple cells and the complex cells, respectively, thereby will be of the forms
$\begin{split}r_{\scriptsize\mbox{simple},1}(\theta)&=\frac{\left|\cos\theta\right|}{\sqrt{\cos^{2}\theta+\kappa^{2}\sin^{2}\theta}},\end{split}$ (12)
$\begin{split}r_{\scriptsize\mbox{simple},2}(\theta)&=\frac{\cos^{2}\theta}{\cos^{2}\theta+\kappa^{2}\sin^{2}\theta},\end{split}$ (13)
$\begin{split}r_{\scriptsize\mbox{complex}}(\theta)&=\frac{\left|\cos\theta\right|^{3/2}}{\left(\cos^{2}\theta+\kappa^{2}\sin^{2}\theta\right)^{3/4}},\end{split}$ (14)
with similar angular dependencies within each class for both the purely spatial receptive fields as for the joint spatio-temporal receptive fields.
Figure 4 shows graphs of these orientation selectivity curves, where we can clearly see how the orientation selectivity becomes more narrow for increasing values of the scale parameter ratio $\kappa$, thus establishing a direct link between the elongation and the degree of orientation selectivity for the idealized models of the receptive fields.
Figure 5: Measurements of the orientation tuning of neurons, at different positions in the visual cortex, according to Nauhaus et al. (2008) (Copyright Cell Press with permission). In this figure it can be seen how the orientation tuning changes from broad to sharp, and thus higher degree of orientation selectivity, with increasing distance from the pinwheels, consistent with the qualitative behaviour that would be obtained if the ratio $\kappa$, between the scale parameters in underlying affine Gaussian smoothing step in the idealized models of spatial and spatio-temporal receptive fields, would increase when moving away from the centers of the pinwheels on the cortical surface.
3 Interpretation of the connection between the orientation selectivity and the elongation of receptive fields in relation to biological measurements
In this section, we will compare the results of the above theoretical predictions with biological results concerning the orientation selectivity of visual neurons.
Nauhaus et al. (2008) have measured the orientation tuning of neurons at different positions in the primary visual cortex for monkey and cat. They found that the orientation tuning is broader near the pinwheel centers and sharper in regions of homogeneous orientation preference. Figure 5 shows an overview of their results, where we can see how the degree of orientation selectivity changes rather continuously from broad to sharp with increasing distance from the pinwheel center (from top to bottom in the figure).
In view of our theoretical results, as summarized in Section 2, concerning the orientation selectivity of receptive fields, where the spatial smoothing part is performed based on affine Gaussian kernels, this qualitative behaviour is consistent with what would be the result if the ratio $\kappa$ between the two scale parameters of the underlying affine Gaussian kernels would increase from a lower to a higher value, when moving away from the centers of the pinwheels on the cortical surface. Thus, the presented theory leads to a prediction about a variability in the eccentricity or the elongation of the receptive fields in the primary visual cortex. In the case of pinwheel structures, the behaviour is specifically consistent with a variability in the eccentricity or the elongation of the receptive fields from the centers of the pinwheels towards the periphery.
Let us furthermore consider the theoretical prediction from (Lindeberg 2023), that the shapes of the affine Gaussian derivative-based receptive fields ought to comprise a variability over a larger part of the affine group than mere rotations, to enable affine covariance and (partial) affine invariance at higher levels in the visual hierarchy. Then, if combined with the theoretical orientation selectivity analysis presented in the paper, those predictions are also consistent with the results by Nauhaus et al. (2008), with an additional explanatory power: If the theoretically motivated prediction would hold, then the underlying theoretical model may enable a deeper interpretation of those biological results, in terms of underlying computational mechanisms in the visual receptive fields, to enable specific functional covariance and invariance properties at higher levels in the visual hierarchy.
A highly interesting quantitative measurement to perform, in view of these theoretical results, would hence be to fit parameterized models of the orientation selectivity, according to Equations (12), (13) and (14)
$r_{\lambda}(\theta)=\left(\frac{\left|\cos\theta\right|}{\sqrt{\cos^{2}\theta+\kappa^{2}\sin^{2}\theta}}\right)^{\lambda}$ (15)
to orientation tuning curves of the form recorded by Nauhaus et al. (2008), to get estimates of the distribution of the parameter $\kappa$ over a sufficiently large population of visual neurons, under the assumption that the spatial components of the biological receptive fields can be well modelled by affine Gaussian derivatives. [^1] At the point of writing this article, the author does, however, not have access to the explicit data that would be needed to perform such an analysis.
In (Lindeberg 2023), a theoretical treatment is given concerning covariance properties of visual receptive fields under natural image transformations, specifically geometric image transformations in terms of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations. According to that theory of spatial and spatio-temporal receptive fields, in terms of generalized Gaussian derivatives, the covariance properties of the receptive fields mean that the shapes of the receptive field families should span the degrees of freedom generated by the geometric image transformations. With regard to spatial affine transformations, which beyond spatial scaling transformations do also comprise spatial rotations and non-uniform scaling transformation with different amount of scaling in two orthogonal spatial directions, this theory implies that affine Gaussian kernels ought to be present in the receptive field families corresponding to different values of the ratio $\kappa$ between the spatial scale parameters, to support affine covariance. In (Lindeberg 2023 Section 3.2) suggestions to new biological measurements were further proposed to support (or reject) those hypotheses.
If we would assume that it would be unlikely for the receptive fields to have as strong variability in their orientational selectivity properties as function of the positions of the neurons in relation to the pinwheel structure, as reported in this study, without also having a strong variability in their eccentricity, then by combining the theoretical analysis in this article with the biological results by Nauhaus et al. (2008), that would serve as possible indirect support for the hypothesis concerning an expansion of receptive field shapes over variations in the ratio between the two scale parameters of spatially anisotropic receptive fields. Thus, if we would assume that the biological receptive fields can be well modelled by the generalized Gaussian derivative model based on affine Gaussian receptive fields, then the biological results by Nauhaus et al. (2008) are fully consistent with the prediction of such an explicit expansion over shapes of the visual receptive fields, based on the orientation selectivity of visual receptive fields, whose spatial smoothing component can be well modelled by affine Gaussian kernels.
Based on these results we propose that, beyond an expansion over rotations, as is performed in current models of the pinwheel structure of visual receptive fields (Bonhoeffer and Grinvald 1991; Blasdel 1992; Swindale 1996; Petitot 2003; Baspinar et al. 2018; Liu and Robinson 2022), also an explicit expansion over the eccentricity $\epsilon$ of the receptive fields (the inverse of the parameter $\kappa$) should be included, when modelling the pinwheel structure in the visual cortex.
Possible ways, by which an explicit dependency on the eccentricity of the receptive fields could be incorporated into the modelling of pinwheel structures, will be outlined in more detail in the following treatment regarding more specific biological hypotheses.
Figure 6: Orientation map in the primary visual cortex of cat, as recorded by Koch et al. (2016) (OpenAccess), with the orientation preference of the receptive fields encoded in terms of colours, and demonstrating that the visual cortex performs an explicit expansion of the receptive field shapes over spatial image orientations. A working hypothesis in the paper concerns to investigate whether the primary visual cortex could additionally perform an expansion over the eccentricity or the elongation of the spatial components of the receptive fields. One possible way of performing such an additional expansion, is over the spatial covariance matrices $\Sigma$ of the affine Gaussian derivative kernels according to (1), and as illustrated in Figure 7, although that illustration would additionally need to be complemented by an identification of opposite image orientations, as, for example, can be achieved by a mapping to the double angle $\varphi\mapsto 2\varphi$, as well as a possible adjustment of the second degree of freedom, in how the variability of the two scale parameters in the affine Gaussian derivative model, beyond a variability over their ratio, varies from the center to the periphery in that illustration.
First-order affine Gaussian derivative kernels
Figure 6: Orientation map in the primary visual cortex of cat, as recorded by Koch et al. (2016) (OpenAccess), with the orientation preference of the receptive fields encoded in terms of colours, and demonstrating that the visual cortex performs an explicit expansion of the receptive field shapes over spatial image orientations. A working hypothesis in the paper concerns to investigate whether the primary visual cortex could additionally perform an expansion over the eccentricity or the elongation of the spatial components of the receptive fields. One possible way of performing such an additional expansion, is over the spatial covariance matrices $\Sigma$ of the affine Gaussian derivative kernels according to (1), and as illustrated in Figure 7, although that illustration would additionally need to be complemented by an identification of opposite image orientations, as, for example, can be achieved by a mapping to the double angle $\varphi\mapsto 2\varphi$, as well as a possible adjustment of the second degree of freedom, in how the variability of the two scale parameters in the affine Gaussian derivative model, beyond a variability over their ratio, varies from the center to the periphery in that illustration.
Figure 7: Distribution of first-order affine Gaussian derivative kernels of the form (1) for different spatial covariance matrices $\Sigma$, with their elements parameterized according to $C_{11}=\sigma_{1}^{2}\,\cos^{2}\varphi+\sigma_{2}^{2}\,\sin^{2}\varphi$, $C_{12}=C_{21}=(\sigma_{1}^{2}-\sigma_{2}^{2})\cos\varphi\,\sin\varphi$, and $C_{22}=\sigma_{1}^{2}\,\sin^{2}\varphi+\sigma_{2}^{2}\,\cos^{2}\varphi$, with the larger spatial scale parameter $\sigma_{2}$ in this illustration held constant, while the smaller scale parameter $\sigma_{1}$ varies as $\sigma_{1}=\sigma_{2}/\kappa$, according to a distribution on a hemisphere. The spatial directional derivatives are, in turn, defined according to $\partial_{\varphi}=\cos\varphi\,\partial_{x_{1}}+\sin\varphi\,\partial_{x_{2}}$. The possible additional variability of the scale parameters, beyond their ratio $\kappa$, is, however, not explicitly addressed in this paper. From a biological viewpoint, one could, indeed, possibly think that it might be easier to keep the smaller scale parameter $\sigma_{1}$ constant, and let the larger scale parameter $\sigma_{2}$ increase towards the periphery, since a higher degree of orientation selectivity can then be achieved by just integrating over successively larger support regions in the image space. The important aspect of this illustration is rather that the eccentricity $\kappa$ increases from the most isotropic image position towards the periphery. With regard to the possible connection to the pinwheel structure in the primary visual cortex, the center in this figure would correspond to the center of the pinwheel, whereas the periphery would correspond to the boundaries of the part of the visual cortex that is closest to the center of one particular pinwheel. (Reprinted from Lindeberg (2023) (OpenAccess).)
4 Explicit testable hypotheses for biological experiments
Based on the above theoretical analysis, with its associated theoretical predictions, we propose that it would be highly interesting to perform experimental characterization and analysis based on joint estimation of (i) orientational selectivity, (ii) receptive field eccentricity, (iii) orientational homogeneity and (iv) location of the neuron in the visual cortex in relation to the pinwheel structure, in the primary visual cortex of animals with clear pinwheel structures, to determine if there is a variability in the eccentricity or elongation of the receptive fields, and specifically if the degree of elongation increases with the distance from the centres of the pinwheels towards periphery, as arising as one possible interpretation of combining the theoretical results about orientation selectivity of affine Gaussian receptive fields in this article with the biological results by Nauhaus et al. (2008).
If additionally, reconstructions of the receptive field shapes could be performed for the receptive fields probed during such a systematic investigation of the difference in response characteristics with the distance from the pinwheel centers, and if the receptive fields could additionally be reasonably well modelled according to the generalized Gaussian model for receptive fields studied and used in this paper, it would be interesting to investigate if the shapes of the affine Gaussian components of these receptive fields would span a larger part of the affine group, than the span over mere image orientations, as already established in the orientation maps of the visual cortex, as characterized by Bonhoeffer and Grinvald 1991, Blasdel 1992 and others, see Figure 7 for an illustration.
If we would lay out the shapes of affine Gaussian receptive fields according to the shapes of their underlying spatial covariance matrices $\Sigma$, then we would for a fixed value of their size (the spatial scale parameter) obtain a distribution of the form shown in Figure 7. That directional distribution is, however, in a certain aspect redundant, since opposite orientations on the unit circle are represented by two explicit copies, where the corresponding receptive fields are either equal, for receptive fields corresponding to spatial directional derivatives of even order, or of opposite sign for derivatives of even order. Could it be established that the receptive fields shapes, if expanded over a variability over eccentricity or elongation, for animals that have a clear pinwheel structure, have a spatial distribution that can somehow be related to such an idealized distribution, if we collapse opposite image orientations to the same image orientation, by e.g. a double-angle mapping $\varphi\mapsto 2\varphi$?
Notably the variability of the spatial covariance matrices in the affine Gaussian derivative model comprises a variability over two spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$, while the theoretical analysis of the orientation selectivity properties studied in this article has mainly concerned their ratio $\kappa=\sigma_{2}/\sigma_{1}$. Hence, the illustration in Figure 7 should not be taken as a literal prediction, even if reduced by a double-angle representation. In Figure 7, the larger scale parameter $\sigma_{1}$ is held constant, for convenience of graphical illustration, as obtained by mapping the uniformly sized receptive fields from a uniform distribution on the hemisphere. More generally, one could also conceive other distributions as possible, such as, for example, instead keeping the smaller eigenvalue $\sigma_{2}$ constant from the center towards the periphery.
To conclude, we propose to state the following testable hypotheses for biological experiments:
Hypothesis 1 (Variability in eccentricity) Let $\sigma_{\varphi}$ and $\sigma_{\bot\varphi}$ be the characteristic lengths in the preferred directions of an orientation selective simple cell in the primary visual cortex. Then, over a population of such simple cells, there is a substantial variability in their eccentricity ratio $\epsilon=\sigma_{\varphi}/\sigma_{\bot\varphi}$.
Hypothesis 2 (Variability in eccentricity coupled to orientational homogeneity) Assuming that Hypothesis 1 holds, let $\epsilon$ denote the eccentricity of a simple cell in the primary visual cortex, and let $H$ be a measure of the homogeneity in the orientation preference of its surrounding neurons. Then, over a population of simple cells, there is a systematic connection between $\epsilon$ and $H$.
Hypothesis 3 (Variability in eccentricity coupled to the pinwheel structure) Assuming that Hypothesis 1 holds, let $\epsilon$ denote the eccentricity of a simple cell in the primary visual cortex. Then, over a population of simple cells, there is a systematic connection between $\epsilon$ and the distance from the nearest pinwheel center.
If Hypothesis 3 would hold, then we could also sharpen this hypothesis further as:
Hypothesis 4 (Increase in elongation with increasing distance from the centers of the pinwheels) Assuming that Hypothesis 3 holds, let $\epsilon$ denote the eccentricity measure of a simple cell in the primary visual cortex defined such that $\epsilon=1$ if the characteristic lengths of the spatial receptive fields are equal, and tending towards zero as the characteristic lengths differ more and more. Then, over a population of simple cells, the eccentricity measure decreases from the center of the pinwheel towards the periphery.
Note that the latter explicit hypotheses have been expressed on a general form, of not explicitly assuming that the biological receptive fields can be well modelled according to the generalized Gaussian derivative model for receptive fields. The essential factor in the definitions is only that it should be possible to define estimates of the characteristic lengths $\sigma_{\varphi}$ and $\sigma_{\bot\varphi}$, so as to be able to define a measure of the eccentricity $\epsilon$.
If either Hypothesis 2 or Hypotesis 3 would hold, then we could also explicitly state the following hypothesis:
Hypothesis 5 (Pinwheel structure more structured than a mere expansion over spatial orientations) The pinwheel structure comprises an, at least, two-dimensional variability of receptive field shapes, beyond an expansion over spatial orientations, also an expansion over the eccentricity of the receptive fields in the primary visual cortex.
For simplicity, we have above expressed these hypotheses for the case of simple cells, for which it is easiest to define the measures $\sigma_{\varphi}$ and $\sigma_{\bot\varphi}$ of the characteristic lengths, because of the linearity of the receptive fields. Provided that corresponding measures of characteristic lengths could also be in a sufficiently well-established way be defined also for non-linear complex cells, corresponding explicit hypotheses could also be formulated for complex cells.
It should finally be stressed that, we have in this treatment not considered the binocular aspects of the pinwheel structure. In Hypothesis 5, the variability of the pinwheel structure over contributions from the left and the right eyes should therefore not be counted as a property to contribute to the terminology “more structured”.
5 Quantitative measurements for detailed characterization
To further characterize possible relationships between the orientational selectivity, receptive field eccentricity, orientational homogeneity, and the location of the neuron in relation to the pinwheel structure in the primary visual cortex, we would also propose to characterize the possible relationships between these entities in terms of:
Quantitative measurement 1: (Relationship between orientational selectivity and receptive field eccentricity) Graph or scatter diagram showing how a quantitative measure of orientational selectivity is related to a quantitative measure of receptive field eccentricity, accumulated over a sufficiently large population of neurons.
Quantitative measurement 2: (Relationship between orientational homogeneity and receptive field eccentricity) Graph or scatter diagram showing how a quantitative measure of orientational homogeneity is related to a quantitative measure receptive field eccentricity, accumulated over a sufficiently large population of neurons.
Quantitative measurement 3: (Relationship between receptive field eccentricity and the pinwheel structure) Graph or scatter diagram showing how a quantitative measure of receptive field eccentricity depends on the distance to the nearest pinwheel center, accumulated over a sufficiently large population of neurons.
Quantitative measurement 4: (Relationship between receptive field eccentricity and the pinwheel structure) Two-dimensional map showing how a quantitative measure of receptive field eccentricity relates to a two-dimensional map of the orientation preference over the same region in the primary visual cortex, with the center of the pinwheel structure explicitly marked, again accumulated over a sufficiently large population of neurons.
If the above theoretically motivated biological hypotheses could be investigated experimentally, and if the above quantitative measurements of receptive field characteristics could be performed, it could be judged if the prediction from the presented theoretical analysis about a systematic variability in receptive field eccentricity, with a possible relationship to the pinwheel structure, could be either experimentally supported or rejected. In a corresponding manner, such a judgement could also answer if the receptive fields in the primary visual cortex could be regarded as spanning a larger part of the affine group, than an expansion over mere rotations in the image domain.
6 Relations to previous work
Beyond the work by Nauhaus et al. (2008), which the treatment in this paper largely builds upon, there is a large body of work on characterizing the orientation selectivity of neurons, by Rose and Blakemore (1974), Schiller et al. (1976), Ringach et al. (2002), Nauhaus et al. (2008), Scholl et al. (2013), Sadeh and Rotter (2014), Goris et al. (2015) and Sasaki et al. (2015), as well as concerning biological mechanisms for achieving orientation selectivity by Somers et al. 1995, Sompolinsky and Shapley 1997, Carandini and Ringach 1997, Lampl et al. 2001, Ferster and Miller 2000, Shapley et al. 2003, Seriès et al. 2004 and Priebe (2016). The focus of this paper work is, however, not on the neural mechanisms that lead to orientation selectivity, but concerning purely functional properties at the macroscopic level.
Mathematical models of biological receptive fields have beyond in terms of Gaussian derivatives (Koenderink and van Doorn 1984; 1987; 1992; Young and his co-workers 1987; 2001; 2001, Lindeberg 2013a; 2021) also been formulated in terms of Gabor filters (Marcelja 1980; Jones and Palmer 1987a; 1987b; Porat and Zeevi 1988). Gaussian derivatives have, in turn, been used as primitives in theoretical models of early visual processing by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Wallis and Georgeson (2009), Hansen and Neumann (2008), Wang and Spratling (2016) and Pei et al. (2016).
Hubel and Wiesel (1959; 1962; 2005) pioneered the study of simple and complex cells. The properties of simple cells have been further characterized by DeAngelis et al. (1995; 2004), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008) and De and Horwitz (2021), and the properties of complex cells investigated by Movshon et al. (1978), Emerson et al. (1987), Touryan et al. (2002; 2005), Rust et al. (2005) and Goris et al. (2015), as well as modelled computationally by Adelson and Bergen (1985), Heeger (1992), Serre and Riesenhuber (2004), Einhäuser et al. (2002), Kording et al. (2004), Merolla and Boahen (2004), Berkes and Wiscott (2005), Carandini (2006) and Hansard and Horaud (2011). In this work, we have followed a specific way of modelling simple and complex cells in terms of affine Gaussian derivatives, according to the generalized affine Gaussian derivative model for visual receptive fields.
7 Summary and discussion
Based on a combination of theoretical results, concerning a direct connection between the degree of orientation selectivity and the degree of elongation of receptive fields, in combination with the results from biological experiments by Nauhaus et al. (2008), which show a substantial variability in the degree of orientation selectivity of the neurons in the primary visual cortex, we have addressed the question of whether the receptive fields in the primary visual cortex can be regarding as spanning a variability in the degree of elongation.
An in-depth theoretical analysis in a companion paper (Lindeberg 2024), as summarized in Section 2, shows how the orientation selectivity depends on a scale parameter ratio $\kappa$, between the scale parameters in the image orientations perpendicular to vs. parallel with the preferred orientation of the receptive fields, for the generalized Gaussian derivative model for visual receptive fields. The explicit expressions for the resulting orientation tuning curves, derived from closed-form theoretical analysis, show that, for both (i) models of simple cells in terms of first-order affine Gaussian derivatives, (ii) models of simple cells in terms of second-order affine Gaussian derivatives, and (iii) models of complex cells in terms of quasi-quadrature measures of simple cells in terms of first- and second-order Gaussian derivatives, the orientation selectivity becomes more narrow with increasing values of the scale parameter ratio $\kappa$. This result holds for both purely spatial models of the receptive fields as well as for joint spatio-temporal receptive field models.
By comparisons with previously established results by Nauhaus et al. (2008), we have in Section 3 demonstrated that the predictions from our theory are consistent with general behaviour of visual neurons in the primary visual cortex, in that the orientation tuning is broader near the pinwheels, while the orientation selectivity becomes sharper further away.
By combining these two results, if we, would take the liberty of assuming that it would be unlikely for the population of receptive fields to exhibit a strong variability in the orientational selectivity of the receptive fields, without also similarly exhibiting a coupled variability in the eccentricity or the the degree of elongation of the receptive fields, then we would have possible indirect support for a previously formulated biological hypothesis in (Lindeberg 2023), stating that the family of receptive field shapes should span the degrees of freedom in the natural geometric image transformations. That way of assumption-based reasoning would then specifically imply indirect support for the hypothesis that the receptive field shapes should span a sufficiently wide range of ratios between the scale parameters in the directions perpendicular to vs. parallel with the preferred orientation of the receptive field, to support affine covariance of the family of visual receptive fields.
Without explicitly relying on expressing such an explicit assumption, regarding whether the visual receptive fields in the primary visual cortex could be well modelled by affine Gaussian derivative based receptive fields, we can, however, firmly state that the biological measurements performed by Nauhaus et al. (2008) are, in combination with the theoretical results summarized in Section 2, consistent with the hypothesis that the receptive fields should span a variability in the eccentricity of the receptive fields. In this respect, the measurements that demonstrate a strong variability in orientation selectivity would specifically be consistent with the theoretically based hypothesis formulated in (Lindeberg 2023), that the receptive fields in the primary visual cortex should span the variability of receptive field shapes under spatial affine transformations.
If we would apply a similar type of assumption-based logical reasoning to the pinwheel structure in the primary visual cortex, then such a reasoning, based on the results by Nauhaus et al. (2008), that the orientation selectivity appears to vary strongly from the centers of the pinwheels towards the periphery, would imply that the pinwheel structure in the visual cortex would, beyond an explicit expansion over image orientations, also comprise an explicit expansion over the eccentricity or the degree of elongation of the receptive fields. Based on these predictions, we have specifically proposed that explicit dependencies on a variability in the eccentricity of the receptive fields should be included, when modelling the pinwheel structures in the primary visual cortex.
Strictly, and formally, the results from such logical inference could, however, only be regarded as theoretical predictions, to generate explicit hypothesis concerning the distribution of receptive field characteristics in these respects. To raise the question of determining if these theoretical predictions would hold in reality, we have in Section 4 formulated a set of testable explicit biological hypotheses, that could be either verified or rejected in neurophysiological experiments, concerning possible variabilities in the eccentricity of the receptive fields in the primary visual cortex of higher mammals, as well as hypotheses about possible connections between such variabilities in the eccentricity or the elongation and other receptive field characteristics, in particular in relation to the pinwheel structure in the primary visual cortex of higher mammals. We have also in Section 5 formulated a set of quantitative measurements that could be made, to characterize how a possible variability in the eccentricity of the receptive field could be related to other receptive field characteristics, including the pinwheel structure in the visual cortex.
Concerning possible limitations in the hypothetical reasoning stages used for for possible logical inference and for formulating the explicit biological hypotheses above, based on explicitly stated assumptions regarding whether the biological receptive fields could be reasonably well modelled by affine Gaussian derivative based receptive fields, to be able to draw possible further conclusions, the possible validity of those hypothetical logical reasoning stages could, however, break down, if there would be other external factors, no covered by the theoretical model, that could also strongly influence the orientation selectivity of the receptive fields. The possible applicability of the hypothetical logical reasoning stages above, does, thus, ultimately depend on possible agreement between the model and biological data, and can only be taken further by performing more detailed actual model fitting (not performed here, because of lack of access to the data by Nauhaus et al. (2008) as well as lack of access to data with receptive field recordings over a sufficiently large population of visual neurons in the primary visual cortex) and/or performing complementary neurophysiological experiments, to ultimately judge if the theoretically based predictions, stated more explicitly in Section 4, would hold for actual biological neurons.
In addition, we have also with the presented treatment demonstrated how the generalized Gaussian derivative model for visual receptive fields lends itself to closed form theoretical analysis, that can lead to well-defined predictions between the orientation selectivity and the anisotropy or elongation of receptive fields in the primary visual cortex, and in this way qualitatively very reasonable predictions regarding phenomena in biological vision.
We have specifically used these results to express possible indirect support for a biological hypothesis concerning a possible expansion of receptive field shapes in the primary visual cortex over the degrees of freedom in natural geometric image transformations, regarding an expansion over the spatial eccentricity (the inverse of the scale parameter ratio $\kappa$) of the receptive fields. Based on these results, we specifically propose to also include such an expansion over eccentricity in models of the pinwheel structure of the receptive fields in the visual cortex. More generally, to ultimately judge if these theoretically based predictions would hold in the primary visual cortex of higher mammals, we have formulated a set of explicit hypotheses and quantitative measurements in Sections 4 and 5, that could serve as a guide to future neurophysiological experiments.